OpenStudy (anonymous):

A projectile is thrown at an angle 30 degree with the horizontal and Kinetic Energy 'E' calculate the Energy in terms of E at maximum height?

OpenStudy (anonymous):

at max height the velocity is always zero. right ?

OpenStudy (anonymous):

Vertical component will be 0

OpenStudy (anonymous):

yah !

OpenStudy (anonymous):

Hmax = u^2 / 8g P.E = mu^2/8

OpenStudy (anonymous):

At hmax ..horizontal component is u cos 30 shuld we have to consider it ? @experimentX

OpenStudy (jfraser):

yes, you still have to consider the horizontal velocity of the projectile. It's KE can't be zero at max height.

OpenStudy (anonymous):

So at Hmax the Energy is same as the initial Energy right?

OpenStudy (jfraser):

no. the energy at max height must be less than the energy at the start. E at max height will be equal to Ecos30

OpenStudy (anonymous):

Hw.......?

OpenStudy (anonymous):

At max point the body will have P.E and K.E due to horizontal component right?

OpenStudy (anonymous):

P Hmax = u^2 / 8g P.E = mu^2/8

OpenStudy (jfraser):

PE will be due to height, KE will be due to its VELOCITY. the only velocity at max height is horizontal velocity. The horizontal velocity doesn't change from the moment of launch.

OpenStudy (anonymous):

E = P.E + 1/2 m u^2 cos^2 30

OpenStudy (jfraser):

yes

OpenStudy (anonymous):

mu^2/8 + 3mu^2/8 = mu^2 /2 which is same as initial K.E ie 1/2 m u^2 cos^2 30 + 1/2 m u^2 sin^2 30 = mu^2/2

OpenStudy (anonymous):

at max height...it'll have only horizontal velocity n horizontal velocity is constant throughout the motion.

OpenStudy (jfraser):

no. your last equation was right.\[E_{total} = PE + KE\]\[E_{total} = mgh + \frac {1}{2} mv^2\] at laungh, PE = 0, so KE = E_total. At max height, PE is also amximum, so the only portion of velocity left is the horizontal\[E_{\max} = E_{total} = mgh + \frac{1}{2}m(vcos30)^2\] @akash123 is corect, the horizontal velocity doesn't change. As I already said

OpenStudy (anonymous):

1/2 m u^2= E

OpenStudy (anonymous):

find u then at max height K.E.= 1/2 m (u cos(theta))^2

OpenStudy (jfraser):

What is it that you really need to know? the TOTAL energy at max height, or the KINETIC energy at max height?

OpenStudy (anonymous):

oh..i am Silly....thxx for reminding @JFraser i need k.E

OpenStudy (anonymous):

K.E at Top most = 3E/4...i got it...)

OpenStudy (anonymous):

1/2 m u^2 3/4 = 3 E/4

OpenStudy (anonymous):

thxxx....guys for helping..)

OpenStudy (vincent-lyon.fr):

I would instinctively say: KEtop = KE horizontal = KE0 cos²30